# Continuum-Field-Theoretical Construction of Conserved Currents, Non-Invertible Symmetries, and Mixed Anomalies in (3+1)D Non-Abelian Topological Order

### v2

@@ VI Summary and Outlook
VI.1 Summary
In this paper, we investigate generalized symmetries in the topological field theory of (3+1)(3+1)D BR topological order. Our strategy is to construct symmetry operators from conserved currents: we extract continuity equations and conserved currents from the equations of motion of the gauge fields and then write down the associated symmetry operators. Geometrically, these generalized symmetries are classified as $1$-form or $2$-form, depending on the dimension of the supporting submanifold. Algebraically, the symmetries include both invertible and non-invertible ones. The invertible higher-form symmetries exhibit a $\mathbb{Z}_{N_{i}}$-type group structure: their symmetry operators obey group-like fusion rules that match the composition law of group elements, as summarized in Sec. IV.3. Equivalently, each such operator admits an inverse whose fusion yields the identity operator. By contrast, the non-invertible symmetry operators (again of either $1$-form or $2$-form type) consist of a unitary operator dressed by a projector. The projector enforces the constraints under which the corresponding current is conserved. Since a projector has no inverse, the resulting symmetry is intrinsically non-invertible. Correspondingly, the fusion of non-invertible symmetry operators can admit multiple fusion channels, as illustrated in Sec. IV.4.
We also analyze anomalies of these generalized symmetries by testing whether they can be gauged. Gauging a single symmetry amounts to coupling its current to a background field in a gauge-invariant way. However, when two symmetries are gauged simultaneously, a mixed anomaly may arise: the action with both coupling terms fails to remain gauge invariant. In some cases, such a mixed anomaly can be canceled by anomaly inflow from a higher-dimensional bulk, so that the combined bulk–boundary system is gauge invariant. In other cases, the mixed anomaly is intrinsic and persists because no bulk term can restore gauge invariance; see Table 3.
@@ VI.1 Summary
Overall, our results provide a field-theoretical characterization of invertible and non-invertible higher-form symmetries in 33d topological order. We expect that the current-based construction developed here can be generalized to other 33d topological orders, including those described by quantum doubles of finite groups. In parallel, recent progress has established diagrammatic representations for 33d and 44d topological orders from continuum field theories, yielding stringent consistency conditions for higher-dimensional topological orders [78]. It would be interesting to further clarify how the fusion rules of the symmetry operators constructed here are constrained by, and in turn inform, these diagrammatic consistency conditions.
VI.2 Outlook
@@ Acknowledgements.
This work was in part supported by National Natural Science Foundation of China (NSFC) under Grants No. 12474149 and No. 12274250, Research Center for Magnetoelectric Physics of Guangdong Province under Grant No. 2024B0303390001, and Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices under Grant No. 2022B1212010008.
@@ Appendix C Explanation of no higher-dimensional bulk to cancel mixed anomaly of two non-invertible symmetries
$S_{\text{gauging}}^{m_{1},m_{2}}=S_{aab}+S_{\text{coupling}}.$ (191)
Under $\begin{cases}\mathcal{B}_{m}^{1}\rightarrow&\mathcal{B}_{m}^{1}+dV_{m}^{1}\\ a^{1}\rightarrow&a^{1}-\frac{1}{N_{1}}V_{m}^{1}\end{cases}$ and $\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases}$, the variation of $S_{aab}$ is
$\Delta S_{aab}=$ $\int_{M_{4}}-\frac{1}{2\pi}b^{1}dV_{m}^{1}-\frac{1}{2\pi}b^{2}dV_{m}^{2}+q\left(a^{1}-\frac{1}{N_{1}}V_{m}^{1}\right)\left(a^{2}-\frac{1}{N_{2}}V_{m}^{2}\right)b^{3}-qa^{1}a^{2}b^{3}$
@@
We now simplify $\Delta S_{aab}+\Delta S_{\text{coupling}}$. First, the term $\int_{M_{4}}-\frac{1}{2\pi}b^{1}dV_{m}^{1}-\frac{1}{2\pi}b^{2}dV_{m}^{2}$ from $\Delta S_{aab}$ cancels $\int_{M_{4}}\frac{1}{2\pi}dV_{m}^{1}b^{1}+\frac{1}{2\pi}dV_{m}^{2}b^{2}$ from $\Delta S_{\text{coupling}}$. Assuming $M_{4}$ is closed, we use
@@
$\int_{M_{4}}\underbrace{-\frac{1}{2}\frac{1}{2\pi}\frac{N_{2}}{2\pi}dV_{m}^{1}a^{2}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}-\underbrace{\frac{1}{2}\frac{1}{2\pi}V_{m}^{1}\frac{N_{2}}{2\pi}a^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{1}}V_{m}^{1}a^{2}b^{3}$ from $\Delta S_{aab}$}}=0.$ (195)
@@
$\int_{M_{4}}\underbrace{\frac{1}{2}\frac{1}{2\pi}dV_{m}^{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}-\underbrace{\frac{1}{2}\frac{1}{2\pi}\frac{N_{1}}{2\pi}a^{1}V_{m}^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{2}}a^{1}V_{m}^{2}b^{3}$ from $\Delta S_{aab}$}}=0.$ (196)
@@
$\int_{M_{4}}\underbrace{\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{1}V_{m}^{2}\beta^{3}-\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{2}V_{m}^{1}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}+\underbrace{\frac{1}{2}\frac{1}{4\pi^{2}}V_{m}^{1}V_{m}^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{1}N_{2}}V_{m}^{1}V_{m}^{2}b^{3}$ from $\Delta S_{aab}$ }}=0.$ (197)
Up to this point, we have reduced $\Delta S_{aab}+\Delta S_{\text{coupling}}$ to
$\Delta S_{aab}+\Delta S_{\text{coupling}}=$ $\int_{M_{4}}-\frac{1}{2}\frac{q}{N_{1}}V_{m}^{1}a^{2}b^{3}-\frac{1}{2}\frac{q}{N_{2}}a^{1}V_{m}^{2}b^{3}+\frac{1}{2}\frac{q}{N_{1}N_{2}}V_{m}^{1}V_{m}^{2}b^{3}$
@@
We can now explain why we cannot find a bulk term $S_{\text{bulk}}$ such that $\Delta S_{aab}+\Delta S_{\text{coupling}}+\Delta S_{\text{bulk}}=0$ under the gauge transformations $\begin{cases}\mathcal{B}_{m}^{1}\rightarrow&\mathcal{B}_{m}^{1}+dV_{m}^{1}\\ a^{1}\rightarrow&a^{1}-\frac{1}{N_{1}}V_{m}^{1}\end{cases}$ and $\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases}$. The action $S_{\text{bulk}}$ must be a functional of background fields only, since $a^{i}$ and $b^{i}$ live only in (3+1)(3+1)D. However, $\Delta S_{aab}+\Delta S_{\text{coupling}}$ contains terms that depend explicitly on $a^{i}$ and $b^{i}$. Such terms cannot be canceled by $\Delta S_{\text{bulk}}$, because $\Delta S_{\text{bulk}}$ cannot involve $a^{i}$ or $b^{i}$. Therefore, a bulk term $S_{\text{bulk}}$ satisfying $\Delta S_{aab}+\Delta S_{\text{coupling}}+\Delta S_{\text{bulk}}=0$ does not exist.

### v3

@@ VI Summary and Outlook
In this work we presented a constructive, current-to-defect route to generalized symmetries—including intrinsically non-invertible higher-form symmetries—in a class of (3+1)(3+1)D twisted $BF$ topological field theories with an $a\wedge a\wedge b$ twist and gauge group $G=\prod_{i}\mathbb{Z}_{N_{i}}$. These TQFTs serve as effective continuum descriptions of Borromean-Rings (BR) topological order, i.e., three-dimensional non-Abelian topological orders supporting Borromean-Rings braiding. The starting point is the field-theory counterpart of “symmetry $\Leftrightarrow$ conserved current”: we derive continuity equations directly from the equations of motion, identify the corresponding conserved charges, and exponentiate them to obtain topological symmetry operators/defects supported on codimension-$p$ submanifolds. In this way, the generalized-symmetry data are obtained in a step-by-step and computable manner directly from the continuum action: currents $\rightarrow$ defects $\rightarrow$ fusion/anomaly diagnostics.
A central structural result is that the equations of motion naturally organize the conserved currents into two qualitatively distinct classes, which in turn lead to two different symmetry structures:
  • Type-I currents are conserved identically (in the sense of Bianchi/Noether-type identities) and generate invertible higher-form symmetries. Their symmetry operators admit inverses and obey group-like fusion, reproducing the expected higher-form composition laws (with minimal generators of $\mathbb{Z}_{N_{i}}$ type); see Sec. IV.3.
  • Type-II currents become conserved only after imposing additional consistency conditions on admissible gauge-field configurations. At the operator/defect level, these conditions are implemented by projector dressings of otherwise unitary topological operators. This provides an explicit and practical mechanism for intrinsic non-invertibility in (3+1)(3+1)D: the resulting higher-form symmetry operators do not admit inverses, and their fusion is generically multi-channeled, with the channel structure dictated by the fusion of the projector factors; see Sec. IV.4.
With explicit (non-)invertible symmetry operators in hand, we then computed their fusion algebra by composing defects supported on the same submanifold. This yields a unified picture in which group-like fusion for invertible symmetries and multi-channel fusion for non-invertible symmetries arise from the same current-based construction, with the projector constraints providing a computable organizing principle for the non-group-like sector.
We further diagnosed anomalies by coupling conserved currents to background gauge fields and testing gaugeability. While each generalized symmetry can be coupled in a gauge-invariant manner individually, simultaneous coupling of two symmetries can exhibit a mixed anomaly, signaled by an obstruction to maintaining gauge invariance. We identify two qualitatively distinct outcomes: (i) mixed anomalies that can be canceled by anomaly inflow from a one-higher-dimensional topological field theory, and (ii) intrinsic gauging obstructions already encoded in the (3+1)(3+1)D continuum theory, for which no bulk counterterm restores gauge invariance; see Table 3.
@@
Overall, our results provide a systematic field-theoretic characterization of invertible and non-invertible higher-form symmetries and their mixed anomalies in (3+1)(3+1)D TQFTs, organized as a practical pipeline from equations of motion to currents, symmetry defects, fusion, and gaugeability. We expect that this construction can be extended to other families of three-dimensional non-Abelian topological orders (including quantum doubles of finite groups) and can serve as a useful interface between condensed-matter realizations and the defect/anomaly language natural in SymTFT. In parallel, recent progress has established diagrammatic representations and consistency conditions for 33d and 44d topological orders from continuum field theories [76]. It would be interesting to clarify how the fusion rules derived here are constrained by, and in turn sharpen, these diagrammatic consistency conditions.
@@ Acknowledgements.
This work was in part supported by National Natural Science Foundation of China (NSFC) under Grants No. 12474149 and No. 12274250.
@@ Appendix B An example of fusion rule in the setup of $G=\left(\mathbb{Z}_{6}\right)^{3}$ and $p=2$
This means that
$\mathbb{P}_{123}\times\mathbb{P}_{234}=$ $\frac{1}{6}\times\frac{1}{3}\times\frac{1}{3}\sum_{l_{1}=0}^{5}\sum_{l_{2},l_{3}=0}^{2}e^{{\rm i}2l_{1}\int_{\gamma_{p}}a^{1}+{\rm i}l_{2}\cdot 2\int_{\gamma_{p}}a^{2}+{\rm i}l_{3}\cdot 2\int_{\sigma}b^{3}}.$ (188)
@@
This means that
$\mathbb{P}_{351}\times\frac{1}{3}\sum_{m_{2}=0}^{2}e^{{\rm i}m_{2}\cdot 2\int_{\gamma_{p}}a^{2}}=$ $\underbrace{\frac{1}{6}\sum_{h_{2}=0}^{5}e^{{\rm i}2h_{2}\int_{\gamma_{p}}a^{1}}\times\frac{1}{6}\sum_{h_{3}=0}^{5}e^{{\rm i}4h_{3}\cdot\left(\int_{\sigma}b^{3}+\int_{\gamma_{p}}a^{2}\right)}}_{\mathbb{P}_{351}}\times\frac{1}{3}\sum_{m_{2}=0}^{2}e^{{\rm i}m_{2}\cdot 2\int_{\gamma_{p}}a^{2}}$
$=$ $\frac{1}{6}\sum_{h_{2}=0}^{5}e^{{\rm i}2h_{2}\int_{\gamma_{p}}a^{1}}\times\frac{1}{3}\sum_{h_{3}=0}^{2}e^{{\rm i}h_{3}\cdot\left(2\int_{\sigma}b^{3}\right)}\times\frac{1}{3}\sum_{m_{2}=0}^{2}e^{{\rm i}m_{2}\cdot 2\int_{\gamma_{p}}a^{2}}$
$=$ $\mathbb{P}_{123}\times\mathbb{P}_{234}$ (190)
@@ Appendix C Explanation of no higher-dimensional bulk to cancel mixed anomaly of two non-invertible symmetries
$S_{\text{gauging}}^{m_{1},m_{2}}=S_{\text{BR}}+S_{\text{coupling}}.$ (193)
Under $\begin{cases}\mathcal{B}_{m}^{1}\rightarrow&\mathcal{B}_{m}^{1}+dV_{m}^{1}\\ a^{1}\rightarrow&a^{1}-\frac{1}{N_{1}}V_{m}^{1}\end{cases}$ and $\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases}$, the variation of $S_{\text{BR}}$ is
$\Delta S_{\text{BR}}=$ $\int_{M_{4}}-\frac{1}{2\pi}b^{1}dV_{m}^{1}-\frac{1}{2\pi}b^{2}dV_{m}^{2}+q\left(a^{1}-\frac{1}{N_{1}}V_{m}^{1}\right)\left(a^{2}-\frac{1}{N_{2}}V_{m}^{2}\right)b^{3}-qa^{1}a^{2}b^{3}$
@@
We now simplify $\Delta S_{\text{BR}}+\Delta S_{\text{coupling}}$. First, the term $\int_{M_{4}}-\frac{1}{2\pi}b^{1}dV_{m}^{1}-\frac{1}{2\pi}b^{2}dV_{m}^{2}$ from $\Delta S_{\text{BR}}$ cancels $\int_{M_{4}}\frac{1}{2\pi}dV_{m}^{1}b^{1}+\frac{1}{2\pi}dV_{m}^{2}b^{2}$ from $\Delta S_{\text{coupling}}$. Assuming $M_{4}$ is closed, we use
@@
$\int_{M_{4}}\underbrace{-\frac{1}{2}\frac{1}{2\pi}\frac{N_{2}}{2\pi}dV_{m}^{1}a^{2}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}-\underbrace{\frac{1}{2}\frac{1}{2\pi}V_{m}^{1}\frac{N_{2}}{2\pi}a^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{1}}V_{m}^{1}a^{2}b^{3}$ from $\Delta S_{\text{BR}}$}}=0.$ (197)
@@
$\int_{M_{4}}\underbrace{\frac{1}{2}\frac{1}{2\pi}dV_{m}^{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}-\underbrace{\frac{1}{2}\frac{1}{2\pi}\frac{N_{1}}{2\pi}a^{1}V_{m}^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{2}}a^{1}V_{m}^{2}b^{3}$ from $\Delta S_{\text{BR}}$}}=0.$ (198)
@@
$\int_{M_{4}}\underbrace{\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{1}V_{m}^{2}\beta^{3}-\frac{1}{2}\frac{1}{4\pi^{2}}dV_{m}^{2}V_{m}^{1}\beta^{3}}_{\text{from $\Delta S_{\text{coupling}}$}}+\underbrace{\frac{1}{2}\frac{1}{4\pi^{2}}V_{m}^{1}V_{m}^{2}\frac{pN_{3}}{N_{123}}b^{3}}_{\text{$=\frac{1}{2}\frac{q}{N_{1}N_{2}}V_{m}^{1}V_{m}^{2}b^{3}$ from $\Delta S_{\text{BR}}$ }}=0.$ (199)
Up to this point, we have reduced $\Delta S_{\text{BR}}+\Delta S_{\text{coupling}}$ to
$\Delta S_{\text{BR}}+\Delta S_{\text{coupling}}=$ $\int_{M_{4}}-\frac{1}{2}\frac{q}{N_{1}}V_{m}^{1}a^{2}b^{3}-\frac{1}{2}\frac{q}{N_{2}}a^{1}V_{m}^{2}b^{3}+\frac{1}{2}\frac{q}{N_{1}N_{2}}V_{m}^{1}V_{m}^{2}b^{3}$
@@
We can now explain why we cannot find a bulk term $S_{\text{bulk}}$ such that $\Delta S_{\text{BR}}+\Delta S_{\text{coupling}}+\Delta S_{\text{bulk}}=0$ under the gauge transformations $\begin{cases}\mathcal{B}_{m}^{1}\rightarrow&\mathcal{B}_{m}^{1}+dV_{m}^{1}\\ a^{1}\rightarrow&a^{1}-\frac{1}{N_{1}}V_{m}^{1}\end{cases}$ and $\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases}$. The action $S_{\text{bulk}}$ must be a functional of background fields only, since $a^{i}$ and $b^{i}$ live only in (3+1)(3+1)D. However, $\Delta S_{\text{BR}}+\Delta S_{\text{coupling}}$ contains terms that depend explicitly on $a^{i}$ and $b^{i}$. Such terms cannot be canceled by $\Delta S_{\text{bulk}}$, because $\Delta S_{\text{bulk}}$ cannot involve $a^{i}$ or $b^{i}$. Therefore, a bulk term $S_{\text{bulk}}$ satisfying $\Delta S_{\text{BR}}+\Delta S_{\text{coupling}}+\Delta S_{\text{bulk}}=0$ does not exist.